OpenStudy (anonymous):

Find the factors of 2a2 + 9a − 18. (2a − 3)(a + 6) (2a + 3)(a − 6) (2a − 9)(a + 2) (2a − 2)(a + 9)

OpenStudy (akashdeepdeb):

2a2 + 9a − 18. I can give you the answer directly but do you want to know the ACTUAL method?

OpenStudy (anonymous):

C

OpenStudy (akashdeepdeb):

Not quite! :) Do you want any explanation? :)

OpenStudy (anonymous):

yes please

OpenStudy (akashdeepdeb):

Let this be a quadratic eqqation! ax2 + bx + c = 0 Now to find the roots we have to do something known as splitting the middle term like I had done in the previos video [Only difference in that case a was = 1 and here it is 2] So to find out the middle terms we need 2 digits that multiply to a*c AND add or subtract to 'b' [Like in the last case they were 8 and 3 because when subtracted they gave b which was 5 and when multiplied they gave a*c = 24]

OpenStudy (anonymous):

still dont get it

OpenStudy (akashdeepdeb):

So in this case we need to find out numbers that multiply to 36 and add/subtract to give 9! Two such numbers are? 12 and 3 [12-3 = 9 and 12*3 = 36] Understood? :) So write it like this! 2a2 + 9a − 18. 2a2 + 12a - 3a - 18 2a(a+6) - 3(a+6) (2a-3)(a+6) Got this? :)

OpenStudy (anonymous):

yes thank you

OpenStudy (akashdeepdeb):

:)